OpenStudy (notgoodgirl):

what are the formula for mean, median and mode that equals to 6?

OpenStudy (hedgepig):

seems like you're missing some info

OpenStudy (notgoodgirl):

This is only to help me in one answers

OpenStudy (notgoodgirl):

*question

OpenStudy (hedgepig):

is this everything you're given?

OpenStudy (notgoodgirl):

I only want to know that

OpenStudy (usukidoll):

x.x this is like a 6 6 6 thing think about it 6 6 6 6 mode is repeated numbers which is 6 the middle of 6 6 6 6 is 6 the mean 6+6+6+6/4 =24/4=6

OpenStudy (notgoodgirl):

ok thanks